# Evidence-backed Fact Checking using RAG and Few-Shot In-Context Learning with LLMs

### v1

@@ Abstract
Given the widespread dissemination of misinformation on social media, implementing fact-checking mechanisms for online claims is essential. Manually verifying every claim is highly challenging, underscoring the need for an automated fact-checking system. This paper presents our system designed to address this issue. We utilize the Averitec dataset Schlichtkrull et al. (2023) to assess the veracity of claims. In addition to veracity prediction, our system provides supporting evidence, which is extracted from the dataset. We develop a Retrieve and Generate (RAG) pipeline to extract relevant evidence sentences from a knowledge base, which are then inputted along with the claim into a large language model (LLM) for classification. We also evaluate the few-shot In-Context Learning (ICL) capabilities of multiple LLMs. Our system achieves an ’Averitec’ score of 0.33, which is a 22% absolute improvement over the baseline. All code will be made available on All code will be made available on https://github.com/ronit-singhal/evidence-backed-fact-checking-using-rag-and-few-shot-in-context-learning-with-llms.
@@ 1 Introduction
In this paper, we present our system for automated fact verification. Our system classifies a given textual claim into one of four categories: Supported, Refuted, Conflicting Evidence/Cherrypicking, or Not Enough Evidence. Additionally, it provides supporting evidence for the classification. Our approach leverages recent advancements in Large Language Models (LLMs), specifically Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL), to produce evidence-backed veracity predictions. Given a claim and a collection of documents, our system first employs a RAG pipeline to retrieve the three most relevant documents and extract evidence from them. Subsequently, we utilize ICL to determine the veracity of the claim based on the extracted evidence. We evaluated our system on the Averitec dataset Schlichtkrull et al. (2023), where it outperforms the official baseline by a large margin. Our key contributions are as follows:
@@ 2 Related Work
Recently, there has been increased research interest in fake news detection and fact checking. Glazkova et al. (2021) proposed and ensemble of BERT Devlin et al. (2019) for Covid fake news Patwa et al. (2021b) detection. Harrag and Djahli (2022) employ deep learning techniques for fact checking in Arabic Baly et al. (2018). Song et al. (2021) tackle the problems of fake news detection using graph neural networks. The factify tasks Mishra et al. (2022); Suryavardan et al. (2023b) aim to detect multi-modal fake news. However, these systems only provide the veracity prediction without any evidence.
On the fever dataset Thorne et al. (2018), Krishna et al. (2022) design a seq2seq model to generate natural logic-based inferences as proofs, resulting SoTA performance on the dataset. Schuster et al. (2021) release the VitaminC dataset and propose contrastive learning for fact verification. Hu et al. (2022) propose a DRQA retriever Chen et al. (2017) based method for fact checking over unstructured information Aly et al. (2021). These systems provide evidence or explanation to back their predictions but they test the veracity of synthetic claims whereas we test real claims.
Some researchers have also used LLMs to tackle the problem. Kim et al. (2024) leverage multiple LLMs as agents to enhance the faithfulness of explanations of evidence for fact-checking. Zhang and Gao (2023) design a hierarchical prompting method which directs LLMs to separate a claim into several smaller claims and then verify each of them progressively.
There have also been attempts to solve the problem using RAG. Khaliq et al. (2024) utilize multimodal LLMs with a reasoning method called chain of RAG to provide evidence based on text and image. Deng et al. (2024) propose a method to decrease misinformation in RAG pipelines by re-ranking the documents during retrieval based on a credibility score assigned to them. Similar to these systems, we also use RAG and LLMs in our solution.
@@ 3 Data
We utilize the Averitec dataset Schlichtkrull et al. (2023) for fact-checking purposes. This dataset comprises claims accompanied by a knowledge store (a collection of articles). Each claim is annotated with question-answer pairs that represent the evidence, a veracity label, and a justification for the label. The veracity label can be one of the following: Support (S), Refute (R), Conflicting Evidence/Cherrypicking (C), or Not Enough Evidence (N). A claim is labeled as C when it contains both supporting and refuting evidence. The data distribution, as shown in Table 1, indicates a class imbalance favoring the R class, while the C and N classes have relatively few examples. The final testing was conducted on 2,215 instances. For further details on the dataset, please refer to Schlichtkrull et al. (2023).
On average, each claim contains 17 words. Figure 1 presents a word cloud of the claims, revealing that most claims pertain to topics related to politics and COVID-19.
Figure 1: Word cloud of the claims. We can see that Politics and COVID-19 are common topics among the claims.
@@ 4 Methodology
Claim:
Donald Trump has stated he will not contest for the next elections
Figure 3: The prompt used for generating questions. Some correct and incorrect examples are given to guide the LLM.
@@ 4.1 Document Retrieval Using Dense Embeddings
Once the claim is embedded, we utilize Facebook’s FAISS (Facebook AI Similarity Search) Douze et al. (2024) to conduct a nearest-neighbor search within the knowledge store. FAISS is an efficient library for similarity search and clustering of dense vectors. We configured FAISS to retrieve the top three documents most relevant to the claim. These documents are then used in the subsequent evidence extraction and veracity prediction steps.
@@ 4.2 Evidence Extraction Using LLMs
Question Generation: The claim is transformed into a question that challenges its validity using an LLM. We employ In-Context Learning, which enables the model to generate responses based on a few provided examples, aiding in the creation of nuanced and contextually appropriate questions. The prompt is designed to ensure that the generated question challenges the claim’s veracity rather than simply seeking a factual answer. An example prompt is provided in Figure 3.
Answer Generation: After generating the question, we provide a single document to a LLM and pose the question. The LLM is prompted to deliver concise and definitive answers derived directly from the content of the document. An example of this prompt is illustrated in Figure Y. This process is repeated for each of the three documents, resulting in three distinct answers for each claim. These answers collectively constitute our evidence. It is important to note that, in our experiments, the LLM used for answer generation does not necessarily need to be the same as the one used for question generation. The prompt utilized in this step is similar to the one depicted in Figure 2.
@@ 4.3 Few-Shot ICL for Final Classification
For the final veracity prediction, we use an LLM to classify a claim based on the three pieces of evidence extracted earlier. The LLM is prompted to choose one out of the four possible classes. The prompt is designed to guide the model through the classification process, ensuring that it correctly interpreted the relationship between the claim and the evidence. An example prompt is given in figure 4.
Our methodology aligns with recent advancements in retrieval-augmented generation (RAG) models which alleviate hallucination and ICL methods, which have been shown to improve the accuracy of LLMs. The integration of these state-of-the-art methods is an attempt to ensure that the extracted evidence is both relevant, contextually appropriate and for validating the claims accurately.
@@ 5 Experiments
To convert documents into dense embeddings, we utilize the dunzhang/stella_en_1.5B_v5 model [^1] https://huggingface.co/dunzhang/stella_en_1.5B_v5. This model was chosen because, at the time of our experiments, it was ranked first on the Massive Text Embedding Benchmark (MTEB) leaderboard Muennighoff et al. (2022), and it currently holds the second position.
@@
For answer generation and final classification, we experiment with multiple LLMs of varying sizes, including InternLM2.5 Cai et al. (2024), Llama-3.1 Dubey et al. (2024), Phi-3-medium Abdin et al. (2024), Qwen2 Yang et al. (2024), and Mixtral Jiang et al. (2024). These models were selected based on their performance on the Open LLM Leaderboard Fourrier et al. (2024) and their availability through Ollama.
We utilize an A40 GPU for Mixtral, while all other models are run on an A100 GPU. Our best-performing model, Mixtral, requires an average of 2 minutes for evidence extraction and final prediction. All code will be made available on https://github.com/ronit-singhal/evidence-backed-fact-checking-using-rag-and-few-shot-in-context-learning-with-llms.
@@ 6 Results and Analysis
Table 4: Class-wise F1 scores of our top three LLMs on the dev set. Classes are Supported (S), Refuted (R), Not enough evidence (N) and conflicting evidence/cherrypicking. Macro-averaged F1 score is also reported.
@@
Given that Mixtral achieved the highest Averitec score on the development set, we selected it for evaluation on the test set. Table 3 provides a comparison of our system and the official baseline Schlichtkrull et al. (2023) on the test set. The baseline model utilizes Bloom Scao et al. (2023) for evidence generation, followed by re-ranking of the evidence using a a finetuned BERT-large model and finally a a finetuned BERT-large model veracity prediction. Unlike the baseline, which uses finetuned models, we only use a few train examples via ICL. Despite that, our system outperforms the baseline across all three evaluation metrics. Notably, our Averitec score of 0.33 is a 22% absolute improvement over the baseline.
6.1 Class-wise Performance
@@ 7 Conclusion and Future Work
In this paper, we have introduced our system for evidence-supported automated fact verification. Our system – based on RAG and ICL – requires only a minimal number of training examples to extract relevant evidence and make veracity predictions. We observed that all LLMs demonstrate sub-optimal performance on the "Conflicting Evidence/Cherrypicking" and "Not Enough Evidence" categories, which emphasizes the inherent challenges of these categories. Additionally, no single LLM consistently outperforms others across all categories. Our system achieved an Averitec score of 0.33, highlighting the complexity of the problem and indicating a substantial potential for future improvement
@@ 8 Limitation
As we are using few-shot ICL, our system cannot make use of large annotated datasets if available, because of the limitation of the prompt size. Furthermore, we assume the availability of high quality LLMs, which might not be the case for some low-resource languages.

### v2

@@ Abstract
Given the widespread dissemination of misinformation on social media, implementing fact-checking mechanisms for online claims is essential. Manually verifying every claim is very challenging, underscoring the need for an automated fact-checking system. This paper presents our system designed to address this issue. We utilize the Averitec dataset Schlichtkrull et al. (2023) to assess the performance of our fact-checking system. In addition to veracity prediction, our system provides supporting evidence, which is extracted from the dataset. We develop a Retrieve and Generate (RAG) pipeline to extract relevant evidence sentences from a knowledge base, which are then inputted along with the claim into a large language model (LLM) for classification. We also evaluate the few-shot In-Context Learning (ICL) capabilities of multiple LLMs. Our system achieves an ’Averitec’ score of 0.33, which is a 22% absolute improvement over the baseline. Our Code is publicly available on https://github.com/ronit-singhal/evidence-backed-fact-checking-using-rag-and-few-shot-in-context-learning-with-llms.
@@ 1 Introduction
Figure 1: Overview diagram of our system. First, the claim is used to retrieve the top 3 relevant documents. Next, evidence is extracted from these documents using questions and answers generated by an LLM. Finally, the evidence is used for veracity prediction via few-shot ICL.
In this paper, we present our system for automated fact verification. Our system classifies a given textual claim into one of four categories: Supported, Refuted, Conflicting Evidence/Cherrypicking, or Not Enough Evidence. Additionally, it provides supporting evidence for the classification. Our approach leverages recent advancements in Large Language Models (LLMs), specifically Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL), to produce evidence-backed veracity predictions. Given a claim and a collection of documents, our system first employs a RAG pipeline to retrieve the three most relevant documents and extract evidence from them. Subsequently, we utilize ICL to determine the veracity of the claim based on the extracted evidence. Figure 1 provides a high-level overview of our system. We evaluate our system on the Averitec dataset Schlichtkrull et al. (2023), where it outperforms the official baseline by a large margin. Our key contributions are as follows:
@@ 2 Related Work
Recently, there has been increased research interest in fake news detection and fact checking. Glazkova et al. (2021) proposed an ensemble of BERT Devlin et al. (2019) for Covid fake news Patwa et al. (2021b) detection. Harrag and Djahli (2022) employed deep learning techniques for fact checking in Arabic Baly et al. (2018). Song et al. (2021) tackled the problem of fake news detection using graph neural networks. The factify tasks Mishra et al. (2022); Suryavardan et al. (2023b) aimed to detect multi-modal fake news. However, these systems only provide the veracity prediction without any evidence.
On the FEVER dataset Thorne et al. (2018), Krishna et al. (2022) designed a seq2seq model to generate natural logic-based inferences as proofs, resulting in SoTA performance on the dataset. Schuster et al. (2021) released the VitaminC dataset and propose contrastive learning for fact verification. Hu et al. (2022) proposed a DRQA retriever Chen et al. (2017) based method for fact checking over unstructured information Aly et al. (2021). These systems provide evidence or explanation to back their predictions but they test the veracity of synthetic claims whereas we test real claims.
Some researchers have also used LLMs to tackle the problem. Kim et al. (2024) leveraged multiple LLMs as agents to enhance the faithfulness of explanations of evidence for fact-checking. Zhang and Gao (2023) designed a hierarchical prompting method which directs LLMs to separate a claim into several smaller claims and then verify each of them progressively.
There have also been attempts to solve the problem using RAG. Khaliq et al. (2024) utilized multimodal LLMs with a reasoning method called chain of RAG to provide evidence based on text and image. Deng et al. (2024) proposed a method to decrease misinformation in RAG pipelines by re-ranking the documents during retrieval based on a credibility score assigned to them. Similar to these systems, we also use RAG and LLMs in our solution.
@@ 3 Data
Figure 2: Word cloud of the claims. We can see that Politics and COVID-19 are common topics in the claims.
We utilize the Averitec dataset Schlichtkrull et al. (2023) for fact-checking purposes. This dataset comprises claims accompanied by a knowledge store (a collection of articles). Each claim is annotated with question-answer pairs that represent the evidence, a veracity label, and a justification for the label. The veracity label can be one of the following: Support (S), Refute (R), Conflicting Evidence/Cherrypicking (C), or Not Enough Evidence (N). A claim is labeled as C when it contains both supporting and refuting evidence. The data distribution, as shown in Table 1, indicates a class imbalance favoring the R class, while the C and N classes have relatively few examples. The final testing is conducted on 2,215 instances Schlichtkrull et al. (2024). For further details on the dataset, please refer to Schlichtkrull et al. (2023, 2024).
On average, each claim consists of 17 words. Figure 2 (word cloud of the claims) reveals that most claims are related to politics and COVID-19.
@@ 4 Methodology
Claim:
Donald Trump has stated he will not contest for the next elections
Figure 3: The prompt used for generating questions. Some manually created correct and incorrect examples are given to guide the LLM.
@@ 4.1 Document Retrieval Using Dense Embeddings
Once the claim is embedded, we utilize FAISS (Facebook AI Similarity Search) Douze et al. (2024) to conduct a nearest-neighbor search within the knowledge store. FAISS is an efficient library for similarity search and clustering of dense vectors. We configure FAISS to retrieve the top three documents most relevant to the claim. These documents are then used in the subsequent evidence extraction and veracity prediction steps.
@@ 4.2 Evidence Extraction Using LLMs
Question Generation: The claim is transformed into a question challenging its validity using an LLM. We employ In-Context Learning, which enables the model to generate responses based on a few provided examples, aiding in the creation of nuanced and contextually appropriate questions. The prompt is designed to ensure that the generated question challenges the claim’s veracity rather than simply seeking a factual answer. An example prompt is provided in Figure 3.
Answer Generation: After generating the question, we provide a single document to an LLM and pose the question. The LLM is prompted to deliver concise and definitive answers derived directly from the content of the document. This process is repeated for each of the three documents, resulting in three distinct answers for each claim. These answers collectively constitute our evidence. It is important to note that in our experiments, the LLM used for answer generation does not necessarily need to be the same as the one used for question generation. The prompt utilized in this step is similar to the one depicted in Figure 4.
@@ 4.3 Few-Shot ICL for Final Classification
For the final veracity prediction, we use an LLM to classify a claim based on the three pieces of evidence extracted earlier. The LLM is prompted to choose one out of the four possible classes. The prompt is designed to guide the model through the classification process, ensuring that it correctly interprets the relationship between the claim and the evidence. An example prompt is given in Figure 5.
Our methodology aligns with recent advancements in retrieval-augmented generation (RAG) pipelines which alleviate hallucination and ICL methods, which have been shown to improve the accuracy of LLMs. The integration of these state-of-the-art methods is an attempt to ensure that the extracted evidence is both relevant and contextually appropriate for validating the claims accurately.
@@ 5 Experiments
To convert documents into dense embeddings, we utilize the dunzhang/stella_en_1.5B_v5 model [^1] https://huggingface.co/dunzhang/stella_en_1.5B_v5. This model is chosen because, at the time of our experiments, it was ranked first on the Massive Text Embedding Benchmark (MTEB) leaderboard Muennighoff et al. (2022), and holds the second position at the time of writing this paper.
@@
For answer generation and final classification, we experiment with multiple LLMs of varying sizes, including InternLM2.5 Cai et al. (2024), Llama-3.1 Dubey et al. (2024), Phi-3-medium Abdin et al. (2024), Qwen2 Yang et al. (2024), and Mixtral Jiang et al. (2024). These models are selected based on their performance on the Open LLM Leaderboard Fourrier et al. (2024) and their availability through Ollama.
We utilize an A40 GPU for Mixtral, while all other models are run on an A100 GPU. Our best-performing model, Mixtral, requires an average of 2 minutes for evidence extraction and final prediction. Our code is publicly available on https://github.com/ronit-singhal/evidence-backed-fact-checking-using-rag-and-few-shot-in-context-learning-with-llms.
@@ 6 Results and Analysis
Given that Mixtral achievs the highest Averitec score on the development set, we select it for evaluation on the test set. Table 3 provides a comparison of our system and the official baseline Schlichtkrull et al. (2023) on the test set. The baseline model utilizes Bloom Scao et al. (2023) for evidence generation, followed by re-ranking of the evidence using a finetuned BERT-large model and finally a finetuned BERT-large model veracity prediction. Unlike the baseline, which uses finetuned models, we only use a few train examples via ICL. Despite that, our system outperforms the baseline across all three evaluation metrics. Notably, our Averitec score of 0.33 is a 22% absolute improvement over the baseline.
6.1 Class-wise Performance
@@
Table 4: Class-wise F1 scores of our top three LLMs on the dev set. Classes are Supported (S), Refuted (R), Not enough evidence (N), and conflicting evidence/cherrypicking. Macro-averaged F1 score is also reported.
@@ 7 Conclusion and Future Work
In this paper, we introduced our system for evidence-supported automated fact verification. Our system - based on RAG and ICL - requires only a minimal number of training examples to extract relevant evidence and make veracity predictions. We observed that all LLMs demonstrate sub-optimal performance on the "Conflicting Evidence/Cherrypicking" and "Not Enough Evidence" categories, which emphasizes the inherent challenges of these categories. Additionally, no single LLM consistently outperforms others across all categories. Our system achieved an Averitec score of 0.33, highlighting the complexity of the problem and indicating a substantial potential for future improvement.
@@ 8 Limitation
As we are using few-shot ICL, our system cannot make use of large annotated datasets if available, because of the limitation of the prompt size. Furthermore, we assume the availability of high-quality LLMs, which might not be the case for some low-resource languages.